# Accepted Paper: ISCA 2024, Argentina

### v1

@@ 1. Introduction
Our complete accelerator design is named NeuraChip, a spatial accelerator featuring Coarse-Grained Reconfigurable Array (CGRA) on-chip interconnections. NeuraChip is versatile due to its ability to efficiently compute graphs with varying degrees of sparsity and sparsity patterns.
@@
  • Heterogeneous Processing Approach: We present NeuraChip, an innovative GNN spatial accelerator featuring a decoupled computation pipeline. Segregating multiplication and accumulation operations into dedicated components, we optimize data reuse through strategic mapping.
@@ 2.4. Mapping Algorithms Design
To address these challenges, we propose a novel approach: Dynamic Reseed Hash-Based Mapping (DRHM). This method is similar to prime modular hashing, but with a significant enhancement. After processing a predetermined set of computations, we reseed the hash function. The updated seed values are then stored in a compact lookup table. This dynamic reseeding ensures that the distribution of indices does not become predictable, effectively mimicking the sparsity-agnostic property of random hashing.
@@ 3.3. NeuraCore
The NeuraCore is the primary compute engine in our accelerator. It computes the multiplication operation and generates the partial products of matrix multiplication operations. It is a simple in-order core with support for matrix instructions. NeuraCore supports a special matrix instruction called matrix_mult_hash_4 or simply MMH4.
@@
where $opcode$ represents the operation code, which specifies the MMH4 instruction to be executed by NeuraCore. $Base_{addr}$ denotes the base address used to offset the address of all other addresses involved in this instruction. $A\_data_{addr}$ refers to the memory address where the data of matrix A is located (matrix A is stored in CSC storage format). $B\_col\_ind_{addr}$ points to the memory address containing the column indices of matrix B (matrix B is stored in CSR storage format). $B\_data_{addr}$ indicates the memory address where data from matrix B is stored. $roll\_counter_{addr}$ denotes the memory address where the rolling eviction counter is located. The pseudocode for executing the MMH4 instruction is shown in Algorithm 1. Each MMH4 instruction has the capability to dispatch up to $16$ HACC instructions (further elaborated in the NeuraMem section).
@@ 4. Design Space Exploration of NeuraChip
The flexibility of our NeuraSim simulator, which is used to simulate our NeuraChip accelerator, enables us to evaluate multiple NeuraChip configurations. We have two primary design goals: i) optimizing resource utilization across the accelerator to enhance speedup and ii) striking a balance between performance, chip area, and power consumption to make sure the advantages outweigh the costs (Srinivasan et al., 2016; Banerjee et al., 2004).
@@
Values represent the count of elements per component across three tile configurations.
Tile Size Variation: We introduce three distinct configurations of NeuraChip, named Tile-4, Tile-16, and Tile-64, derived from experimenting with various workloads. The detailed configurations of NeuraCore and NeuraMem components are provided in Table 2, while the overall accelerator configurations for these tile sizes are listed in Table 3. We focus on six key parameters to assess the architectural impact of these configurations, as shown in Figure 11.
Key observations include:
@@
Routers per tile $4$ $16$ $64$
Total Routers $32$ $128$ $512$
@@
Values represent the count of components/elements across the entire NeuraChip accelerators for three different tile configurations.
@@ 5.3. Comparative Analysis with Sparse Matrix Accelerators
As we can see, NeuraChip beats the CPU and GPU computing platforms in all cases. The average performance improvements are $2\times2$ over the CPU, a 17.1×17.1$\times$17.1 × and 13.3×13.3$\times$13.3 × average speedup over the NVIDIA Hopper GPU using the cuSPARSE and CUSP libraries, respectively, and 16.7×16.7$\times$16.7 × average speedup over the AMD’s MI100 GPU using hipSPARSE.
@@ Acknowledgements.
This research was supported in part by the Institute for Experiential AI and the NSF IUCRC Center for Hardware and Embedded Systems Security and Trust (CHEST), NSF CNS 2312275, NSF CNS 2312276, and by Samsung Advanced Institute of Technology, Samsung Electronics Co., Ltd. Additionally, we acknowledge the financial assistance from grant RYC2021-031966-I funded by MCIN/AEI/10.13039/501100011033 and the “European Union NextGenerationEU/PRTR.”, and grant PID2022-136315OB-I00 funded by MCIN/AEI/10.13039/501100011033/ and by “ERDF A way of making Europe”, EU.

### v2

@@ 1. Introduction
Our complete accelerator design is named NeuraChip, a spatial accelerator featuring Coarse-Grained Reconfigurable Array (CGRA) on-chip interconnects. NeuraChip is versatile due to its ability to efficiently compute graphs with varying degrees of sparsity and sparsity patterns.
@@
  • Heterogeneous Processing Approach: We present NeuraChip, an innovative GNN spatial accelerator featuring a decoupled computation pipeline (Smith, 1982). Segregating multiplication and accumulation operations into dedicated components, we optimize data reuse through strategic mapping.
@@ 2.4. Mapping Algorithms Design
To address these challenges, we propose a novel approach: Dynamic Reseed Hash-Based Mapping (DRHM). This method is similar to prime modular hashing, but with a significant enhancement. After processing a predetermined set of computations, we reseed the hash function. The updated seed values are then stored in a compact lookup table. This dynamic reseeding ensures that the distribution of indices does not become predictable, effectively mimicking the sparsity-agnostic property of random mapping.
@@ 3.3. NeuraCore
The NeuraCore is the primary compute engine in our accelerator. It computes the multiplication operation and generates corresponding partial products. It is a simple in-order core with support for matrix instructions. NeuraCore supports a special matrix instruction called matrix_mult_hash_4 or simply MMH4.
@@
Algorithm 1 presents the MMH4 instruction execution pseudocode where $opcode$ represents the operation code, which specifies the MMH4 instruction to be executed by NeuraCore. $Base_{addr}$ denotes the base address used to offset the address of all other addresses involved in this instruction. $A\_data_{addr}$ refers to the memory address where the data of matrix A is located (matrix A is stored in CSC storage format). $B\_col\_ind_{addr}$ points to the memory address containing the column indices of matrix B (matrix B is stored in CSR storage format). $B\_data_{addr}$ indicates the memory address where data from matrix B is stored. $roll\_counter_{addr}$ denotes the memory address where the rolling eviction counter is located. The instruction layout for MMH4 is presented in Figure 7. Each MMH4 instruction has the capability to dispatch up to $16$ HACC instructions (further elaborated in the NeuraMem section).
@@ 4. Design Space Exploration of NeuraChip
The flexibility of our NeuraSim simulator enables us to evaluate multiple NeuraChip configurations. We have two primary design goals: i) optimizing resource utilization across the accelerator to enhance speedup and ii) striking a balance between performance, chip area, and power consumption to make sure the advantages outweigh the costs (Srinivasan et al., 2016; Banerjee et al., 2004).
@@
Tile Size Variation: We define a tile as a modular unit that can be configured with varying amounts of computational and memory resources. Each tile is connected to a dedicated HBM memory channel, with NeuraChip hosting a total of eight tiles to match the eight available memory channels. We introduce three distinct configurations of NeuraChip, named Tile-4, Tile-16, and Tile-64, derived from experimenting with various workloads. The detailed configurations of NeuraCore and NeuraMem components are provided in Table 2, while the overall accelerator configurations for these tile sizes are listed in Table 3. We focus on six key parameters to assess the architectural impact of these configurations, as shown in Figure 11. Key observations include:
@@
Routers per tile $4$ $8$ $32$
Total Routers $32$ $64$ $256$
@@ 5.3. Comparative Analysis with Sparse Matrix Accelerators
As we can see, NeuraChip outperforms the CPU and GPU computing platforms in all cases. The average performance improvements are $2\times2$ over the CPU, a 17.1×17.1$\times$17.1 × and 13.3×13.3$\times$13.3 × average speedup over the NVIDIA Hopper GPU using the cuSPARSE and CUSP libraries, respectively, and 16.7×16.7$\times$16.7 × average speedup over the AMD’s MI100 GPU using hipSPARSE.
@@ Acknowledgements.
This research was supported in part by the Institute for Experiential AI and the NSF IUCRC Center for Hardware and Embedded Systems Security and Trust (CHEST), NSF CNS 2312275, NSF CNS 2312276, Samsung Advanced Institute of Technology, Samsung Electronics Co., Ltd., NRF-2023R1A2C2004229, and IITP No.RS-2023-00228255. Additionally, we acknowledge the financial assistance from grant RYC2021-031966-I funded by MCIN/AEI/10.13039/501100011033 and the “European Union NextGenerationEU/PRTR.”, and grant PID2022-136315OB-I00 funded by MCIN/AEI/10.13039/501100011033/ and by “ERDF A way of making Europe”, EU.